OpenStudy (anonymous):

it says simplify each expression 1. 5x^2-x^2 2. 12 divided by (-3/4) 3.-(-2+6t) 4.-3[b-(-7)]

OpenStudy (hba):

Ok So Lets Start With The First One :)

OpenStudy (hba):

Do You Know How To Take Common

OpenStudy (hba):

\[1. 5x^2-x^2 \]

OpenStudy (anonymous):

4x^2

OpenStudy (hba):

Do You Know How To Take Common ? @tanfernan000

OpenStudy (anonymous):

um no haha

OpenStudy (hba):

Ok Do You Know How To Subract ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2. -16

OpenStudy (hba):

Ok Then You Just Subract 5x^2-1x^2 What Do You Get ?

OpenStudy (anonymous):

nooooo clue i havent done this math in 4 years

OpenStudy (hba):

Its Jus 5-1

OpenStudy (hba):

I Hope You Know What 5-1 =

OpenStudy (hba):

@tanfernan000

OpenStudy (anonymous):

4 haha

OpenStudy (anonymous):

@hba Love the way you write, hahahah

OpenStudy (hba):

@tanfernan000 So What Will Be 5x^2-1x^2=

OpenStudy (hba):

@Fall12-13 Thanks :)

OpenStudy (hba):

Dont Worry Jus Answer We Are Here To Help You If You Go Wrong :)

OpenStudy (anonymous):

what about the ^2 ?

OpenStudy (hba):

Ok Tell Me What Will Be 5x-1x=

OpenStudy (anonymous):

4x?!

OpenStudy (hba):

Yeah Right

OpenStudy (hba):

So What Will Be 5x^2-1x^2=

OpenStudy (hba):

@tanfernan000

OpenStudy (anonymous):

4x^2?

OpenStudy (hba):

Yeah Right

OpenStudy (hba):

Good :)

OpenStudy (anonymous):

carry on @hba

OpenStudy (hba):

Lets Look At The Next Question Now

OpenStudy (anonymous):

ok!:)

OpenStudy (hba):

@gsRttN I Think You Should Carry On I Will Be Back

OpenStudy (anonymous):

ok i can try.

OpenStudy (anonymous):

@tanfernan000 are you there

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

@gsRttN

OpenStudy (anonymous):

we will start from 2nd question.. do you know how to divide

OpenStudy (anonymous):

yes i do @gsRttN

OpenStudy (anonymous):

12 divided by (-3/4) for this 12/(-3/4) you can write this as: 12*4/(-3) when 12 is divide by 3 we get 4 . so answer is -4*4 = -16

OpenStudy (hba):

I Am Really Sorry But I Am Back Now

OpenStudy (anonymous):